There are 156 concrete blocks available to make a retaining wall. The bottom three rows will all have the same number of blocks. The next six rows will each have two blocks fewer than the row below it. How many blocks are in each row?
Answer by checkley75(3666)

3x+x-2+x-4+x-6+x-8+x-10+x-12=156
9x-42=156
9x=156+42
9x=198
x=198/9
x=22 is the number of blocks in the first 3 rows.
then you'll have:
20,18,16,14,12 & 10 in each successive row.
proof
3*22+20+18+16+14+12+10=156
66+90=156
156=156